OpenStudy (anonymous):

MEDAL & FAN FOR HELP AND RIGHT ANSWER IM TAKING A QUIZ HURRY! A rectangle has sides of 4 and 8 inches. What is the perimeter of the rectangle, expressed in feet?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

4+4+8+8=24. 2l+2w 2*4+2*8 =? 8+16=24

OpenStudy (anonymous):

24 inches. 12inch=1foot. 2 feet

OpenStudy (anonymous):

24

OpenStudy (anonymous):

soo 24ft?

OpenStudy (anonymous):

no. 24inches. 24 inches into feet will be 2 feet

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

Ft NOT IN

OpenStudy (anonymous):

BUT ITS ASKING FOR FT

OpenStudy (anonymous):

I gave the right answer first. The perimeter is 24 inches but converted into feet it would be 2 feet.

OpenStudy (anonymous):

OOOH

OpenStudy (anonymous):

Medal?

OpenStudy (anonymous):

but it says expressed in feet

OpenStudy (anonymous):

yes.. and i said the perimeter in feet is 2.

OpenStudy (anonymous):

Do you have multiple choice answers?

OpenStudy (anonymous):

no , but I just thought the answer would be in ft and I have to explain myself because im homeschooled